Ty and Jamal are making sandwiches together . Ty can make 3 sandwiches in 2 minutes and 6 in 4 minutes . Jamal can make 1 sandwiches in 2 minutes and 2 in 4 minutes . How many will they have made together in 10 minutes? How many will each have made?
1
Expert's answer
2020-10-12T17:19:31-0400

Ty

"3\/2=6\/4=1.5" sandwiches per minute

"10=4+4+2"

"1.5*10=15" sandwiches

Jamal

"1\/2=2\/4=0.5" sandwiches per minute

"0.5*10=5" sandwiches

Ty + Jamal=15+5=20